OpenStudy (anonymous):

Two men are running in the same direction along a circle. They meet for the first time at a point diametrically opposite the starting point when the faster one is in his fifth round. Find the ratio of their speeds.

OpenStudy (anonymous):



OpenStudy (dumbcow):

do they both start at same point? what round is the slower runner on?

OpenStudy (anonymous):

dont know

OpenStudy (anonymous):

but travel in the same diection but what about starting point

OpenStudy (dumbcow):

ok well since it says they meet for first time it must be the case that the faster runner has just lapped the slower one .... making them on their 4th round therefore in same amount of time, the faster runner has ran 4.5 laps while the other has ran 3.5 laps speed = distance/time ratio = (3.5/t)/(4.5/t) = 3.5/4.5

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

4:5?

OpenStudy (dumbcow):

nope? do you know the answer already?

OpenStudy (anonymous):

only options

OpenStudy (anonymous):

4:5,7:9,6:5

OpenStudy (dumbcow):

does my reasoning make sense? my answer is correct i believe, just not in right form